# SAIP-Net: Enhancing Remote Sensing Image Segmentation via Spectral Adaptive Information Propagation

### v1

@@ 3.3. Composite Dilated Convolution Layer
(11) $\mathbf{y}_{i}=\operatorname{Conv}_{r_{i}}(\mathbf{x}_{i}),\quad i\in\{1\ldots D\}.$
@@
(12) $\mathcal{F}_{\text{CDC}}(Y^{l})=\operatorname{Conv}\Big{(}\Big{[}y_{1}\ldots y_{D}\Big{]}\Big{)}.$
@@
3.4. Learnable High-Pass Filter Stem
@@
Figure 5. (a). Overview of the proposed LhpfStem. The module is composed of a stack of Lhpf layers. In a Lhpf layer, the high-pass output is computed by subtracting the low-pass response features from the input, thereby enhancing edge details and fine structures. (b). Our module extracts key high-frequency features, sharpens feature boundaries, and improves segmentation performance.
Figure 6. Visual comparison of segmentation results on the Potsdam and LoveDA dataset. The regions enclosed by pink ellipses indicate areas with improved intra-class consistency, and the areas marked by orange ellipses highlight enhanced boundary accuracy.
@@
(13) $\tilde{W}(p,q)=\frac{\operatorname{Softmax}(W_{\text{l}})(p,q)\cdot H(p,q)}{\sum_{p^{\prime},q^{\prime}\in\Omega}\bar{W}(p^{\prime},q^{\prime})\cdot H(p^{\prime},q^{\prime})},$
@@
(14) $\tilde{X}(i,j)=\sum_{p,q\in\Omega}X(i+p,j+q)\,\tilde{W}(p,q).$
@@
(15) $\text{HighPass}(x)(i,j)=X(i,j)-\tilde{X}(i,j).$
@@
3.5. Loss Functions
To train SAIP-Net, we adopt a hybrid loss that combines Cross-Entropy (CE) and Dice(Milletari et al., 2016) losses, balancing pixel-wise classification and region-level overlap. The total loss is defined as:
(16) $\mathcal{L}_{\text{total}}=\lambda\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(\hat{y}_{i,c})\right)+(1-\lambda)\left(1-\frac{2\sum_{i,c}y_{i,c}\hat{y}_{i,c}}{\sum_{i,c}y_{i,c}+\sum_{i,c}\hat{y}_{i,c}}\right),$
@@ 3.5. Loss Functions
4. Experiments
4.1. Implementation Details
@@
To evaluate our model, we conduct experiments on two widely used remote sensing datasets: Potsdam [^1] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-potsdam.aspx and LoveDA(Wang et al., 2021). All results are reported on the test/validation set of LoveDA and the test set of Potsdam, which are strictly held out during training and model selection to avoid data leakage.
@@
Table 1. Comparison of model efficiency and computational cost. SAIP-Net achieves a favorable balance between performance and complexity, with reduced computational cost, particularly in GFLOPs ($1024\times 1024$ input).
@@
4.2. Evaluation Metrics.
We compare our method against several baseline approaches with consistent or larger parameter magnitudes, including UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017), DeepLabV3+ (Chen et al., 2018), DANet (Fu et al., 2019), LANet (Ding et al., 2020), FFPNet (Xu et al., 2020), TransUNet (Chen et al., 2021), UperNet RSP-Swin-T (Wang et al., 2022c), RSSFormer (Xu et al., 2023), FactSeg (Ma et al., 2022), UNetFormer (Wang et al., 2022b), and AerialFormer (Yamazaki et al., 2023). As shown in Table 1, our model achieves comparable or even lower parameter counts and GFLOPs ($1024\times 1024$ input) compared to baseline methods by adopting a more efficient backbone, a lightweight upsampling strategy, an effective receptive field expansion mechanism, and a compact stem with fewer parameters.
@@
4.3. Test Results
Table 2. Quantitative results on the official test set of LoveDA. Bold numbers indicate the best performance in each column.
@@ 4.3. Test Results
Table 3. Quantitative results on the official validation set of LoveDA. Bold numbers indicate the best performance in each column.
Method mIoU $\uparrow$ Bkg. Building Road Water Barren Forest Agri.
UNet 47.33 53.95 56.13 51.26 69.15 19.75 29.00 52.10
TransUNet 50.56 53.18 62.37 52.27 69.06 27.20 34.61 55.19
DeepLabV3+ 50.15 51.75 59.92 53.76 68.35 26.60 39.14 51.53
SegNet 48.34 53.02 54.53 55.42 69.08 16.37 35.67 54.31
AerialFormer-T 52.57 53.80 64.02 55.61 66.90 32.83 45.37 49.46
Our SAIP-Net 54.46 55.36 64.53 57.66 71.77 28.90 43.53 59.44
Table 4. Quantitative results on the Potsdam dataset. Bold numbers indicate the best performance in each column.
@@
The left part of Figure 6 presents segmentation results on the LoveDA dataset, which includes high-resolution images of both urban and rural scenes with complex semantics. Our method is evaluated on both the validation and test sets of LoveDA, with detailed quantitative results provided in Table 3 and Table 2, respectively. The results demonstrate that our model consistently achieves competitive performance across different classes. Visual comparisons further highlight the robustness of our approach in handling diverse object appearances and fine-grained details inherent to remote sensing imagery.
The Potsdam dataset features diverse urban scenes captured under various conditions. As summarized in Table 4, our method achieves the highest mIoU, OA, and mF1 scores, demonstrating its capability to handle the complexities of urban remote sensing imagery. Notably, as illustrated in the right part of Figure 6, our approach shows superior performance in capturing fine segmentation details and accurately delineating class boundaries.
@@
4.4. Ablation Study
Figure 7. Qualitative comparison of ablation variants. The regions enclosed by pink ellipses indicate areas with improved intra-class consistency, and the areas marked by orange ellipses highlight enhanced boundary accuracy.
@@ 4.4. Ablation Study
Table 5. Ablation Study Results on the Potsdam dataset and LoveDA validation set. We use DeepLabV3+ as the baseline model, progressively replacing its components with our proposed modules. The symbol “✓” indicates the inclusion of the corresponding component, while “–” denotes its absence. T.E indicates transformer encoder.
@@
Table 6. Ablation Study on the Effect of $D$ in CDC Layer on Model Performance. We conducted experiments on the LoveDA validation set and found that setting $D=3$ yields the best performance.
@@
5. Conclusion
In this paper, we introduced SAIP-Net: Enhancing Remote Sensing Image Segmentation via Spectral Adaptive Information Propagation. By adaptively propagating spectral information and enlarging the receptive fields, our approach effectively eliminates disruptive high-frequency information within intra-class regions while preserving essential frequency-domain features along class boundaries. Experimental results confirm that our design significantly improves intra-class consistency and enhances boundary accuracy, thereby substantially improving the segmentation of complex structures in remote sensing images.

### v2

@@ 3.3 Composite Dilated Convolution Layer
$\mathbf{y}_{i}=\operatorname{Conv}_{r_{i}}(\mathbf{x}_{i}),\quad i\in\{1\ldots D\}.$ (11)
@@
$\mathcal{F}_{\text{CDC}}(Y^{l})=\operatorname{Conv}\Big(\Big[y_{1}\ldots y_{D}\Big]\Big).$ (12)
@@
3.4 Learnable High-Pass Filter Stem
@@
Figure 5: (a). Overview of the proposed LhpfStem. The module is composed of a stack of Lhpf layers. In a Lhpf layer, the high-pass output is computed by subtracting the low-pass response features from the input, thereby enhancing edge details and fine structures. (b). Our module extracts key high-frequency features, sharpens feature boundaries, and improves segmentation performance.
@@
$\tilde{W}(p,q)=\frac{\operatorname{Softmax}(W_{\text{l}})(p,q)\cdot H(p,q)}{\sum_{p^{\prime},q^{\prime}\in\Omega}\bar{W}(p^{\prime},q^{\prime})\cdot H(p^{\prime},q^{\prime})},$ (13)
@@
$\tilde{X}(i,j)=\sum_{p,q\in\Omega}X(i+p,j+q)\,\tilde{W}(p,q).$ (14)
@@
$\text{HighPass}(x)(i,j)=X(i,j)-\tilde{X}(i,j).$ (15)
@@
3.5 Loss Functions
To train SAIP-Net, we adopt a hybrid loss that combines Cross-Entropy (CE) and Dice[33] losses, balancing pixel-wise classification and region-level overlap. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\lambda\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(\hat{y}_{i,c})\right)+(1-\lambda)\left(1-\frac{2\sum_{i,c}y_{i,c}\hat{y}_{i,c}}{\sum_{i,c}y_{i,c}+\sum_{i,c}\hat{y}_{i,c}}\right),$ (16)
@@ 3.5 Loss Functions
4 Experiments
4.1 Implementation Details
@@
To evaluate our model, we conduct experiments on two widely used remote sensing datasets: Potsdam [^1] https://www.isprs.org/education/benchmarks/UrbanSemLab/2d-sem-label-potsdam.aspx and LoveDA[34]. All results are reported on the test/validation set of LoveDA and the test set of Potsdam, which are strictly held out during training and model selection to avoid data leakage.
@@
Table 1: Comparison of model efficiency and computational cost. SAIP-Net achieves a favorable balance between performance and complexity, with reduced computational cost, particularly in GFLOPs ($1024\times 1024$ input).
@@
4.2 Evaluation Metrics.
We compare our method against several baseline approaches with consistent or larger parameter magnitudes, including UNet [35], SegNet [36], DeepLabV3+ [11], DANet [37], LANet [38], FFPNet [39], TransUNet [24], UperNet RSP-Swin-T [40], RSSFormer [21], FactSeg [15], UNetFormer [23], [10], LSKNet[43] , DecoupleNet [44] and LOGCAN++ [45]. As shown in Table 1, our model achieves comparable or even lower parameter counts and GFLOPs ($1024\times 1024$ input) compared to baseline methods by adopting a more efficient backbone, a lightweight upsampling strategy, an effective receptive field expansion mechanism, and a compact stem with fewer parameters.
@@
4.3 Test Results
Figure 6: Visual comparison of segmentation results on LoveDA dataset. The regions enclosed by pink ellipses indicate areas with improved intra-class consistency, and the areas marked by orange ellipses highlight enhanced boundary accuracy.
Figure 7: Visual comparison of segmentation results on Potsdam dataset. The regions enclosed by pink ellipses indicate areas with improved intra-class consistency, and the areas marked by orange ellipses highlight enhanced boundary accuracy.
Figure 6 presents segmentation results on the LoveDA dataset, which includes high-resolution images of both urban and rural scenes with complex semantics. Our method is evaluated on both the validation and test sets of LoveDA, with detailed quantitative results provided in Table 2 and Table 3, respectively. The results demonstrate that our model consistently achieves competitive performance across different classes. Visual comparisons further highlight the robustness of our approach in handling diverse object appearances and fine-grained details inherent to remote sensing imagery.
The Potsdam dataset features diverse urban scenes captured under various conditions. As summarized in Table 4, our method achieves the highest mIoU, OA, and mF1 scores, demonstrating its capability to handle the complexities of urban remote sensing imagery. Notably, as illustrated in Figure 7, our approach shows superior performance in capturing fine segmentation details and accurately delineating class boundaries.
Table 2: Quantitative results on the official validation set of LoveDA. Bold numbers indicate the best performance in each column.
Method mIoU $\uparrow$ Bkg. Building Road Water Barren Forest Agri.
UNet 47.33 53.95 56.13 51.26 69.15 19.75 29.00 52.10
TransUNet 50.56 53.18 62.37 52.27 69.06 27.20 34.61 55.19
DeepLabV3+ 50.15 51.75 59.92 53.76 68.35 26.60 39.14 51.53
SegNet 48.34 53.02 54.53 55.42 69.08 16.37 35.67 54.31
AerialFormer-T 52.57 53.80 64.02 55.61 66.90 32.83 45.37 49.46
Our SAIP-Net 54.46 55.36 64.53 57.66 71.77 28.90 43.53 59.44
Table 3: Quantitative results on the official test set of LoveDA. Bold numbers indicate the best performance in each column.
@@ 4.3 Test Results
Table 4: Quantitative results on the Potsdam dataset. Bold numbers indicate the best performance in each column.
@@
4.4 Analysis of Failure Cases
Figure 8: Analysis of failure cases on LoveDA Validation Set. (a). Class-wise normalized confusion matrix, showing that the largest confusion involves the Barren class, which is frequently predicted as agricultural or background. (b). The Barren class (purple) illustrates typical errors such as boundary confusion with nearby man-made edges and “interior erosion” in low-texture regions.
Although SAIP-Net achieves consistent gains on most categories, we observe a noticeable drop for Barren on LoveDA dataset.
Data and appearance factors.
As shown in Figure 8, compared with other categories, Barren in LoveDA is intrinsically ambiguous under RGB due to low texture/contrast and appearance proximity to dry agriculture or background compounded by boundary-level annotation uncertainty. These conditions amplify class ambiguity and increase annotation noise near transitions. Together with the relatively smaller frequency of Barren pixels, the class becomes under-represented and more susceptible to bias during optimization.
Frequency bias induced by our design.
SAIP-Net explicitly enhances high-frequency structures via LhpfStem and HP-guided paths in SAFF. While this helps boundaries, Barren regions are often low-texture, low-contrast. The HP-centric pathway may down-weight interior evidence in such regions, causing “interior erosion”. In addition, the LP branch in SAFF, normalized by softmax, favors smooth responses but is guided by features dominated by man-made edges in mixed scenes; this can pull Barren toward adjacent classes during the second-stage offset-based alignment. Finally, CDC enlarges the receptive field with dilations; when texture is scarce, multi-scale context may inadvertently absorb surrounding semantics, increasing contextual leakage into Barren.
Our design prioritizes boundary accuracy and intra-class consistency at controlled complexity, thus we deliberately avoid class-specific heavy treatments (e.g., large reweighting or post-processing), which leads to a moderate drop on this particular class. We also note a plausible side effect of our spectral priors: high-pass–oriented pathways can slightly erode interiors of textureless regions, and large receptive fields may absorb surrounding context. We view handling such low-texture, look-alike regions as a complementary future direction (e.g., class-aware calibration and mid-band constraints or multimodal cues), aiming to improve Barren without compromising efficiency.
4.5 Ablation Study
Figure 9: Qualitative comparison of ablation variants. The regions enclosed by pink ellipses indicate areas with improved intra-class consistency, and the areas marked by orange ellipses highlight enhanced boundary accuracy.
@@ 4.5 Ablation Study
Table 5: Ablation Study Results on the Potsdam dataset and LoveDA validation set. We use DeepLabV3+ as the baseline model, progressively replacing its components with our proposed modules. The symbol “✓” indicates the inclusion of the corresponding component, while “–” denotes its absence. T.E indicates transformer encoder.
@@
We also ablate the proposed SAFF module to isolate its contribution to robustness. Specifically, we compare the full SAIP-Net against (i) an ablated variant w/o SAFF that removes the module while keeping all other components, training protocol, and hyperparameters unchanged, and (ii) the baseline AerialFormer. To probe robustness under common corruptions, we add synthetic Gaussian noise ($\mathrm{prob}{=}0.5$, $\sigma{=}10$), salt-and-pepper noise ($\mathrm{prob}{=}0.5$, ratio=0.01=0.01), and speckle noise ($\mathrm{prob}{=}0.5$, $\sigma{=}0.1$) at validation time. We report the drop in mIoU relative to the clean setting (lower is better).
As summarized in Table 6, SAIP-Net consistently exhibits the smallest performance degradation across all noise types. Relative to the baseline, SAIP-Net reduces the mIoU drop by $1.87$ (Gaussian), $1.50$ (Speckle), and $2.32$ (S&P). Compared to the w/o SAFF ablation, SAIP-Net further narrows the drop by $0.28$, $0.33$, and $0.74$ mIoU under Gaussian, Speckle, and S&P noise, respectively. These gains align with SAFF’s design goal of suppressing disruptive high-frequency components while preserving task-relevant structures.
Table 6: Performance drop in mIoU ($\downarrow$) under synthetic corruptions at validation time. Values denote the decrease relative to the clean setting (lower is better). We compare SAIP-Net (Ours), its ablation w/o SAFF, and the AerialFormer baseline. Bold numbers indicate the best performance in each row.
Noise Ours w/o SAFF Baseline
Gaussian 0.56 0.84 2.43
Speckle 0.11 0.44 1.61
S&P 2.47 3.21 4.79
@@
Table 7: Ablation Study on the Effect of $D$ in CDC Layer on Model Performance. We conducted experiments on the LoveDA validation set and found that setting $D=3$ yields the best performance.
@@
5 Conclusion
In this paper, we introduced SAIP-Net: Enhancing Remote Sensing Image Segmentation via Spectral Adaptive Information Propagation. By adaptively propagating spectral information and enlarging the receptive fields, our approach effectively eliminates disruptive high-frequency information within intra-class regions while preserving essential frequency-domain features along class boundaries. Experimental results confirm that our design significantly improves intra-class consistency and enhances boundary accuracy, thereby substantially improving the segmentation of complex structures in remote sensing images. A remaining limitation is reduced performance on inherently ambiguous, low-texture categories under RGB-only inputs (e.g., Barren in LoveDA dataset). As future work, we will explore methods to better disambiguate such regions without increasing complexity.